1000cm3 vessel is filled with 2 liquids l1 l2 of relative density of 0.75 and 1.85 respectively. If the relative density of the mixture was found to be 1.5 what volume each liquid enters the mixture


1
Expert's answer
2021-04-06T13:50:34-0400

We have two masses in the vessel. Assume one takes x mL and the other V-x. Then the mass of the mixture in the vessel would be


"M=x\u03c1 _\n1\n\u200b\t\n +(V\u2212x)\u03c1 _\n2"

We know the density of the mixture:

"\\rho=\\frac{M}{V}=\\frac{x\u03c1 _\n1\n\u200b\t\n +(V\u2212x)\u03c1 _\n2}{V}"

"x=V\frac{\rho_2-\rho}{\rho_2-\rho_1}\\x=1000\frac{1.85-1.5}{1.85-0.75}=318\ cm^3\\"

Thus, we have 318 mL of liquid 1 and 682 mL of liquid 2.